OpenStudy (anonymous):

Given the arithmetic sequence an = −3 + 9(n − 1), what is the domain for n?

OpenStudy (campbell_st):

g goes from 1 to infinity... this is because n - 1 > 0 and n is an integer value

OpenStudy (freckles):

Honestly this question makes no sense :(

OpenStudy (campbell_st):

well its for a term in an arthimetic sequence where the 1st term is -3 and the common difference is 9

OpenStudy (freckles):

you could have negative subscripts no this is not likely the answer you could have 0 and positive integer subscripts or you could have positive integer subscripts

OpenStudy (freckles):

but what is preventing us from saying \[a_0=-3+9(0-1) \\ a_0=-3+9(-1)=-3-9=-12\] and called the first term -12 instead @campbell_st or mean we could go with negative subscripts as I said before but they are probably not looking for that they are probably looking for what campbell_st said which is n-1>=0

OpenStudy (campbell_st):

I just wouldn't overthink things a term in an arithmetic sequence is \[A_{n} = a_{0} + (n -1) \times d \] so the get the 1st term \[a_{0} ~~~then ~~~n - 1 = 0 ~~~solve ~for~n\]

OpenStudy (anonymous):

here are the answer choices

OpenStudy (freckles):

there is no right answer because the sequence can be defined over any integer set but your teacher is probably wanting you to say n>=1 where n is integer

OpenStudy (freckles):

and yes >= means greater to or equal to